OpenStudy (anonymous):

State the slope and the y-intercept for the graph of each equation. 1. y = x + 9

OpenStudy (anonymous):

Take the equation of a line in the form y = mx + b where m is the slope and b is the y value of the y-intercept. What do you get for m and b?

OpenStudy (anonymous):

m = 0? and the b is 9 right?

OpenStudy (anonymous):

b is 9, yeah

OpenStudy (anonymous):

but if m were 0 then you would have 0x = 0

OpenStudy (anonymous):

you can look at x as 1x

OpenStudy (anonymous):

oh. I don't remember how to get the m

OpenStudy (anonymous):

would b be (0,9) ? because we have to graph the points too

OpenStudy (anonymous):

yeah--for the y-intercept x will always be 0. good job

OpenStudy (anonymous):

m comes out to 1 because, substituting back into y=mx+b, y = 1*x + 9 gives y= x + 9 which is what we wanted

OpenStudy (anonymous):

so the y-intercept is (0, 9), and the slope is 1

OpenStudy (anonymous):

yay thank you! so m = ( 0, 1)

OpenStudy (anonymous):

I mean 1, 0

OpenStudy (anonymous):

m is just the slope, not a point

OpenStudy (anonymous):

so the slope and m are just 1

OpenStudy (anonymous):

How do i graph it? Do I just go up 1?

OpenStudy (anonymous):

Like up 9 and 1 to the right?

OpenStudy (anonymous):

1 can be rewritten as 1/1, so taking that as rise/run, you go over 1 and up 1, over 1 and up 1, etc. from the point (0, 9)

OpenStudy (anonymous):

ohhh ok! can you help me with the next one?

OpenStudy (anonymous):

Your graph for the first one will look like this

OpenStudy (anonymous):

yep it does :)

OpenStudy (anonymous):

y = 2x - 5

OpenStudy (anonymous):

y = 2x - 5 -2x

OpenStudy (anonymous):

Now Im stuck

OpenStudy (anonymous):

@cholo71796